# Frequency-Domain Characterization of Load Demand from Electrified Highways

### v1

@@ III-A Scenario S1): Perfect Time Synchronization
Admittedly, EVs are unlikely to be perfectly synchronized in practice. EVs may be approximately synchronized under heavy traffic conditions, when drivers may generally follow the 2-second rule. According to this rule, a driver maintains a two-second spacing between their vehicle and the leading vehicle. If the distance $2v$ traveled in two seconds is an integer multiple of $D$, EVs can get synchronized. Holding S1) as a worst-case scenario, we next study some more realistic conditions.
III-B Scenario S2): Independent Uniformly Distributed Timings
@@
A sufficient condition for the PSD of a stochastic signal to exist is that the signal is wide sense stationary (WSS) [20, Ch. 4]. A stochastic signal $p(t)$ is WSS if its mean does not vary with time and its autocorrelation depends only on the lag:
@@
Subsection III-A considered the worst-case scenario where EVs are perfectly synchronized. This subsection examines the opposite extreme of no synchronization between EVs. This scenario, termed S2), can occur when EVs are spaced sufficiently far apart and move freely. Under S2), we model EV timings $t_{n}$ as independent and identically distributed (iid) random variables uniformly drawn from $[0,T)$. It is reasonable to assume that $t_{n}$ is uniformly distributed through any time interval; we select the interval $[0,T)$ without loss of generality for mathematical convenience. We model EV densities $a_{n}$ as independent and identically distributed (iid) random variables with a given mean and variance. It is reasonable to assume that EV densities are independent of EV timings.
@@
Suppose EVs move at constant speed $v$; EV densities $a_{n}$ are iid with mean $\mu_{a}$ and variance $\sigma_{a}^{2}$; and EV timings $t_{n}$ are iid uniformly within $[0,T)$. Then, the DWPT load $p(t)$ is WSS with mean $\mu_{p}(t)=N\mu_{a}c_{0}$ and PSD:
@@
According to Lemma 2, the spectrum of $p(t)$ under S2) contains the same harmonics as in S1). Recall that under S1), the DC term and harmonics of $p(t)$ scale with $N$. Under S2), harmonics do not disappear, but scale differently. To simplify the analysis, suppose $\sigma_{a}^{2}=0$ and $\mu_{a}=a$, so EV densities become deterministic and conform with Corollary 1. Then, Lemma 2 explains that the DC term under S2) scales with $Nac_{0}$, whereas its $m$-th harmonic scales with $\sqrt{N}a|c_{m}|$. The key conclusion from this analysis is that even when EV timings are independent, DWPT harmonics do not disappear, but scale with $\sqrt{N}$ rather than $N$.
III-C Scenario S3): Vehicle Platoons
Vehicles driving on a highway often form platoons, i.e., groups of vehicles moving together at the same speed. This subsection studies the effect of EV platoons on the DWPT signal. Suppose the $N$ EVs served by an ER segment form platoons. To ease the analysis under S3), we assume:
@@ III-C Scenario S3): Vehicle Platoons
    each platoon has the same number of EVs, $Q$;
@@
    the timing for each platoon is iid and uniformly in $[0,T)$;
@@
    EV densities are iid of given mean $\mu_{a}$ and variance $\sigma_{a}^{2}$.
Assumption ii) considers the worst-case scenario where EVs within a platoon move synchronized and at spacings that are integer multiples of $D$. Scenario S3) can be studied by combining S1) and S2). If EVs within a platoon get synchronized, they can be modeled as a single EV of a larger rating. On the other hand, timings across platoons can be modeled as independent and uniformly distributed. Based on these two points, we next derive the PSD of $p(t)$ under S3).
@@
Under the assumptions of S3), the mean of $p(t)$ is $N\mu_{a}c_{0}$ and its PSD is
@@
To quantify the effect of platoon formations, compare Lemma 3 to Lemma 2. To ease the comparison, consider deterministic EV densities by setting $\sigma_{a}=0$ and $\mu_{a}=a$. Although the DC component under S3) remains $Nac_{0}$, harmonics scale now as $\sqrt{NQ}a|c_{m}|$ instead of $\sqrt{N}a|c_{m}|$. Clearly, harmonic components are more prominent in S3) compared to S2). This is expected as platoon formations imply less smoothing across time-shifted replicas of $h_{T}(t)$. Therefore, platoon formations can increase the THC in $p(t)$. When $Q=1$, scenario S3) becomes equivalent to S2). On the other extreme, when $Q=N$, scenario S3) becomes S1).
@@
Our analysis assumes a constant number $N$ of EVs on the ER segment. A more detailed analysis should model $N$ as a random variable. Traffic flow studies usually postulate that EV arrivals follow a Poisson distribution parameterized by the mean arrival rate $\lambda_{a}$, measured in EV/s. If all EVs move at constant speed $v$, the total number $N$ of EVs driving on an ER segment of length $L$ follows a Poisson distribution with mean $\lambda_{N}=\lambda_{a}L/v$; see [21, Sec 8.3]. If $\lambda_{a}=0.3$ EV/s, $L=10$ mi, and $v=55$ mph, we get the mean value of $\lambda_{N}=192$. Moreover, the probability of $N$ being within $\pm 10\%$ of its mean value $\lambda_{N}=192$ is 84%. To simplify the analysis, we treat $N$ as deterministic and set it to its mean value $\lambda_{N}$ determined by $(\lambda_{a},L,v)$.
@@ IV EVs Moving at Constant but Unequal Speeds
We have hitherto assumed that EVs move at a common constant speed, resulting in the DWPT load $p(t)$ being periodic in time. This section considers the scenario where EVs move at constant yet unequal speeds. In this case, although each EV load $p_{n}(t)$ is periodic, the DWPT load $p(t)$ is non-periodic, and thus, has a continuous spectrum.
IV-A Scenario S4): Constant Gaussian-Distributed Speeds
@@
with the FS coefficients defined as in (10). EV densities $a_{n}$ are again modeled as iid with mean $\mu_{a}$ and variance $\sigma_{a}^{2}$. Different from Sec. III, here speeds $v_{n}$ are modeled as iid Gaussian random variables distributed as $v_{n}\sim\mathcal{N}(\mu_{v},\sigma^{2}_{v})$. If $v_{n}$ is Gaussian, the fundamental frequency $\omega_{n}={2\pi v_{n}}/{D}$ is also Gaussian with mean and variance:
@@
Under S4), the total DWPT load can be expressed as
@@
Recall that a linear combination of periodic signals is not necessarily periodic. More precisely, signal $p(t)$ is periodic if and only if all $\omega_{n}$ are integer multiples of a single frequency $\omega_{0}$. Because this condition is practically impossible, we conclude that $p(t)$ is non-periodic and does not inherit properties p1)–p3) discussed in Section III. Even though $p(t)$ does not have an FS expansion, its frequency-domain analysis is still possible if we can determine its PSD. The following lemma establishes that $p(t)$ is WSS and derives its PSD.
@@
Under the assumptions of S4), the total DWPT load $p(t)$ is WSS with mean $\mu_{p}(t)=N\mu_{a}c_{0}$ and PSD
@@
Lemma 4 confirms that the PSD of $p(t)$ under S4) is not discrete, but continuous: It consists of Gaussian bells centered around the frequency harmonics $\pm m\mu_{\omega}$. Consider the $m$-th Gaussian bell. Because its standard deviation scales with $m$, its bandwidth is proportional to $m$ and its peak value is inversely proportional to $m$. Therefore, the Gaussian bells of higher harmonics reduce in amplitude and expand in bandwidth.
If the DWPT spectrum is continuous, we define its THC as:
$\text{THC}_{p}=\sqrt{\frac{2\int_{0}^{\infty}S_{p}(\omega)~d\omega}{\lim_{\epsilon\rightarrow 0}\int_{-\epsilon}^{\epsilon}S_{p}(\omega)~d\omega}}\times 100\%.$ (18)
According to this definition and Lemma 4, the DWPT load has precisely the same THC under S2) and S4). This is because the Gaussian PDF integrates to unity as a function of $\omega$. This observation also implies that the energy contributed by the bell centered around $m\mu_{\omega}$ under S4) equals the energy of harmonic $m\omega_{0}$ under S2). Interestingly, scenario S2) can be derived as the limiting case of S4) when $\sigma_{v}$ tends to zero and $\mu_{v}=v$. This originates from the fact that the Dirac delta function is the limiting case of the Gaussian PDF for vanishing variance.
To ease comparison with previous scenarios, consider again deterministic EV densities by setting $\sigma_{a}=0$ and $\mu_{a}=a$. Then, the DC term under S4) scales with $Nac_{0}$ and the Gaussian bell centered around the $m$-th harmonic scales with $\sqrt{N}a|c_{m}|$. This is consistent because timings have been assumed independent across EVs as in S2).
@@
$\text{\emph{THC}}_{p}$ $=\text{\emph{THC}}_{h}$ $\text{for S1)};$
$\text{\emph{THC}}_{p}$ $=\text{\emph{THC}}_{h}/\sqrt{N}$ $\text{for S2) and S4)};and$
$\text{\emph{THC}}_{p}$ $=\text{\emph{THC}}_{h}\sqrt{Q/N}$ $\text{for S3)}.$
@@ V Periodograms of DWPT Load Demands
This question boils down to whether the periodogram of a finite-duration realization of $p(t)$ relates to its PSD. This is a well-studied topic in statistical signal processing [20]. Let us briefly review some key points. Consider a DWPT load realization $\hat{p}(t)$ spanning the finite interval $t\in[0,T_{p}]$ of duration $T_{p}$. Grid operators typically use windows of 1–5 min to compute spectra of signals in power system studies [16]. The periodogram of $\hat{p}(t)$ is computed as the squared magnitude of its Fourier transform as [20, Sec 4.3]
@@
In the three tests of this section, we computed numerically the periodograms of 1-min realizations $\hat{p}(t)$, generated according to S2)S4). More specifically, we simulated the DWPT load consumed by an $L=16$ km (10 mile)-long ER segment following the specifications of the INDOT ER testbed. EV loads were computed as trapezoidal pulse trains. EV power densities $a_{n}$ were simulated as iid Gaussian random variables with mean $\mu_{a}=94.45$ kW/m and standard deviation $\sigma_{a}=5.55$ kW/m; see [22]. For S2) and S3), speed was set to $v=24.56$ m/s. At this speed, an EV takes about 10 min to cross the ER segment.
@@
Figure 3: Periodograms of 1-min DWPT loads simulated under S2) for two values of $N$. The magnitude and location of harmonics agree with those predicted by Lemma 2. Moreover, as asserted by Lemma 2, the ratio of the two periodograms should be $10\log_{10}(192/64)^{2}=9.54$ dB at DC, and about $10\log_{10}\left({192}/{64}\right)=4.77$ dB at harmonic components.
@@
Periodogram under S2). To compute the periodogram under S2), EV timings were independently uniformly drawn from $[0,T)$, where $T=D/v$. Because independent timings are likely to occur under lighter traffic, we considered two relatively low values of the mean arrival rate $\lambda_{a}$, namely $0.1$ and $0.3$ EV/s, yielding $N=64$ and $192$ EVs, respectively. Figure 3 depicts the periodogram of a single 1-min realization of $\hat{p}(t)$ after applying a Hanning window. This periodogram differs from the PSD of Lemma 2 as it relies on a single, finite-duration realization $\hat{p}(t)$. Nonetheless, the locations of harmonic peaks coincide with those predicted by the analysis. For $N=64$ and $N=192$, the DC component of $\hat{p}(t)$ was $8.70$ MW and $26.13$ MW, respectively. Integrating the two periodograms around the first harmonic yielded $357$ kW and $680$ kW, respectively, for the two values of $N$. This verifies that although the DC component scales with $N$, harmonics scale with $\sqrt{N}$. In addition, we evaluated the THC for multiple realizations $\hat{p}(t)$ and found it to be 3–6% and 1–3% for the two values of $N$, indicating that the THC decreases with $N$.
@@
Figure 4: Periodograms of 1-min DWPT loads simulated under S2)S4) for $N=192$ EVs. Harmonics increase in power under S3) due to platoons. The increase is about $10\log_{10}Q=6.98$ dB per Lemmas 23. The bells under S4) are centered around the harmonics of S2) because $\mu_{v}=55$ mph. Bells reduce in magnitude and spread in bandwidth with $m$ per Lemma 4.
@@
Periodogram under S3). In this case, we simulated $N=192$ EVs moving in platoons of approximately $Q=5$ EVs per platoon. To capture imperfect synchronization, we first sampled the timing of each platoon independently uniformly from $[0,T)$. Then, for each EV within a platoon, we added an extra time shift drawn independently uniformly from $[0,T/5]$. Figure 4 illustrates the obtained periodogram. Although EVs within a platoon are not perfectly synchronized, the obtained spectrum contains higher harmonic content compared to the spectrum under S2). As dictated by Lemmas 2 and 3, although the DC components are identical (barring the deviations of a single periodogram from its mean), harmonics do increase in power under S3), as they now scale as $\sqrt{QN}$ instead of $\sqrt{N}$. In addition, we computed the THC under different realizations and found it to be around $3\%$ under imperfect synchronization and $16\%$ under perfect synchronization within platoons.

### v2

@@ III-A Scenario S1: Perfect Time Synchronization
Admittedly, EVs are unlikely to be perfectly synchronized in practice. EVs may be approximately synchronized under heavy traffic conditions, when drivers may generally follow the 2-second rule. According to this rule, a driver maintains a two-second spacing between their vehicle and the leading vehicle. If the distance $2v$ traveled in two seconds is an integer multiple of $D$, EVs can get synchronized. Holding S1 as a worst-case scenario, we next study some more realistic conditions. Admittedly, EVs are unlikely to be perfectly synchronized in practice. EVs may be approximately synchronized under heavy traffic conditions, when drivers may generally follow the 2-second rule. According to this rule, a driver maintains a two-second spacing between their vehicle and the leading vehicle. If the distance $2v$ traveled in two seconds is an integer multiple of $D$, EVs can get synchronized. Holding S1 as a worst-case scenario, we next study some more realistic conditions.
III-B Scenario S2: Independent Uniformly Distributed Timings
@@
A sufficient condition for the PSD of a stochastic signal to exist is that the signal is wide sense stationary (WSS) [26, Ch. 4]. A stochastic signal $p(t)$ is WSS if its mean does not vary with time and its autocorrelation depends only on the lag:
@@
Section III-A considered the worst-case scenario where EVs are perfectly synchronized. This subsection examines the opposite extreme: no synchronization among EVs. This scenario, termed S2, can occur when EVs are spaced sufficiently far apart and move freely. Under S2, we model EV timings $t_{n}$ as independent and identically distributed (i.i.d.) random variables uniformly drawn from $[0,T)$. It is reasonable to assume that $t_{n}$ is uniformly distributed through any time interval. We select the interval $[0,T)$ without loss of generality for mathematical convenience. We model EV densities $a_{n}$ as i.i.d. random variables of given mean and variance. It is reasonable to assume that EV densities are independent of EV timings.
@@
Suppose EVs move at constant speed $v$; EV densities $a_{n}$ are i.i.d. with mean $\mu_{a}$ and variance $\sigma_{a}^{2}$; and EV timings $t_{n}$ are i.i.d. uniformly within $[0,T)$. Then, the DWPT load $p(t)$ is WSS with mean $\mu_{p}(t)=N\mu_{a}c_{0}$ and PSD:
@@
According to Lemma 2, the spectrum of $p(t)$ under S2 contains the same harmonics as in S1. Recall that under S1, the DC term and harmonics of $p(t)$ scale with $N$. Under S2, harmonics do not disappear, but scale differently. To simplify the analysis, suppose $\sigma_{a}^{2}=0$ and $\mu_{a}=a$, so EV densities become deterministic and conform with Corollary 1. Then, Lemma 2 explains that the DC term under S2 scales with $Nac_{0}$, whereas its $m$-th harmonic scales with $\sqrt{N}a|c_{m}|$. The key conclusion from this analysis is that even when EV timings are independent, DWPT harmonics do not disappear, but scale with $\sqrt{N}$ rather than $N$.
For stochastic signals $p(t)$, the THC in (6) can be expressed in stochastic terms as
$\text{THC}_{p}=\sqrt{\frac{2\int_{0}^{\infty}S_{p}(\omega)~d\omega-\lim_{\epsilon\rightarrow 0}\int_{-\epsilon}^{\epsilon}S_{p}(\omega)~d\omega}{\lim_{\epsilon\rightarrow 0}\int_{-\epsilon}^{\epsilon}S_{p}(\omega)~d\omega}}\times 100\%.$ (17)
For periodic stochastic signals with FS coefficients $\tilde{c}_{m}$, the THC formula of (17) simplifies as
$\text{THC}_{p}=\sqrt{\frac{2\sum_{m=1}^{\infty}\mathbb{E}[|\tilde{c}_{m}|^{2}]}{\mathbb{E}[|\tilde{c}_{0}|^{2}]}},$ (18)
where $\mathbb{E}[|\tilde{c}_{m}|^{2}]$ is the magnitude of PSD $S_{p}(\omega)$ at frequency $m\omega_{0}$. Under the S2 setting of unsynchronized traffic flow, Lemma 2 shows that for the special case of $\sigma_{a}=0$, the THC of the total DWPT load decreases with $\sqrt{N}$ as
$\text{THC}_{p}=\text{THC}_{h}/\sqrt{N}.$
Therefore, the DWPT load may have insignificant frequency content as $N$ grows, as verified numerically in Sections VVII.
III-C Scenario S3: Vehicle Platoons
Vehicles driving on a highway often form platoons, i.e., groups of vehicles moving together at the same speed. This subsection studies the effect of EV platoons on the DWPT signal. Suppose the $N$ EVs served by an ER segment form platoons. To ease the analysis under S3, we assume:
@@ III-C Scenario S3: Vehicle Platoons
    Each platoon has the same number of EVs. The number of EVs per platoon will be denoted by $Q$;
@@
    The timing for each platoon is i.i.d. and uniformly in $[0,T)$;
@@
    EV densities are i.i.d. of given mean $\mu_{a}$ and variance $\sigma_{a}^{2}$.
Assumption ii) considers the worst-case scenario where EVs within a platoon move synchronized and at spacings that are integer multiples of $D$. Scenario S3 can be studied by combining S1 and S2. If EVs within a platoon get synchronized, they can be modeled as a single EV of a larger rating. On the other hand, timings across platoons can be modeled as independent and uniformly distributed. Based on these two points, we next derive the PSD of $p(t)$ under S3.
@@
Under the assumptions of S3, the mean of $p(t)$ is $N\mu_{a}c_{0}$ and its PSD is
@@
To quantify the effect of platoon formations, compare Lemma 3 to Lemma 2. To ease the comparison, consider deterministic EV densities by setting $\sigma_{a}=0$ and $\mu_{a}=a$. Although the DC component under S3 remains $Nac_{0}$, harmonics scale now as $\sqrt{NQ}a|c_{m}|$ instead of $\sqrt{N}a|c_{m}|$. Clearly, harmonic components are more prominent in S3 compared to S2. This is expected as platoon formations imply less smoothing across time-shifted replicas of $h_{T}(t)$. Therefore, platoon formations can increase the THC in $p(t)$. When $Q=1$, scenario S3 becomes equivalent to S2. On the other extreme, when $Q=N$, scenario S3 becomes S1.
@@
Our analysis assumes a constant number $N$ of EVs on the ER segment. A more detailed analysis should model $N$ as a random variable. Traffic flow studies usually postulate that EV arrivals follow a Poisson distribution parameterized by the mean arrival rate $\lambda_{a}$, measured in EV/s. If all EVs move at constant speed $v$, the total number $N$ of EVs driving on an ER segment of length $L$ follows a Poisson distribution with mean $\lambda_{N}=\lambda_{a}L/v$; see [27, Sec 8.3]. If $\lambda_{a}=0.3$ EV/s, $L=10$ mi, and $v=55$ mph, we get the mean value of $\lambda_{N}=192$. Moreover, the probability of $N$ being within $\pm 10\%$ of its mean value $\lambda_{N}=192$ is 84%. To simplify the analysis, we treat $N$ as deterministic and set it to its mean value $\lambda_{N}$ determined by $(\lambda_{a},L,v)$.
@@ IV EVs Moving at Constant but Unequal Speeds
We have hitherto assumed that EVs move at a common constant speed, resulting in the total DWPT load $p(t)$ being periodic. This section considers the scenario where EVs move at constant yet unequal speeds. In this case, although each EV load is periodic, $p(t)$ is aperiodic and thus has a continuous spectrum.
IV-A Scenario S4: Constant Gaussian-Distributed Speeds
@@
with the FS coefficients defined as in (11). EV densities $a_{n}$ are again modeled as i.i.d. with mean $\mu_{a}$ and variance $\sigma_{a}^{2}$. Different from Sec. III, here speeds $v_{n}$ are modeled as i.i.d. Gaussian random variables distributed as $v_{n}\sim\mathcal{N}(\mu_{v},\sigma^{2}_{v})$. If $v_{n}$ is Gaussian, the fundamental frequency $\omega_{n}={2\pi v_{n}}/{D}$ is also Gaussian with mean and variance:
@@
Under S4, the total DWPT load can be expressed as
@@
Recall that a linear combination of periodic signals is not necessarily periodic. More precisely, signal $p(t)$ is periodic if and only if all $\omega_{n}$ are integer multiples of a single frequency $\omega_{0}$. Because this condition is practically impossible, we conclude that $p(t)$ is non-periodic and does not inherit properties p1)–p3) discussed in Sec. III. Even though $p(t)$ does not have an FS expansion, its frequency-domain analysis is still possible if we can determine its PSD. The following lemma establishes that $p(t)$ is WSS and derives its PSD.
@@
Under the assumptions of S4, the total DWPT load $p(t)$ is WSS with mean $\mu_{p}(t)=N\mu_{a}c_{0}$ and PSD
@@
Lemma 4 confirms that the PSD of $p(t)$ under S4 is not discrete, but continuous: It consists of Gaussian-shaped bells or lobes centered around the frequency harmonics $\pm m\mu_{\omega}$. Consider the $m$-th Gaussian lobe. Because its standard deviation scales with $m$, its bandwidth is proportional to $m$, and its peak value is inversely proportional to $m$. Hence, the lobes $\phi(\omega;m\mu_{\omega},m^{2}\sigma_{\omega}^{2})$ reduce in amplitude and expand in bandwidth for higher $m$.
Do the Gaussian lobes overlap? The effective bandwidth of a Gaussian lobe can be shown to be 2.3 times its standard deviation. Here, the effective bandwidth is defined as the difference between two points whose value on the lobe is half the maximum value of the Gaussian lobe. Then, the $m$-th lobe essentially occupies the frequency range:
$\left[m(\mu_{\omega}-1.15\sigma_{\omega}),m(\mu_{\omega}+1.15\sigma_{\omega})\right]$
and the $(m+1)$-th lobe occupies the range:
$\left[(m+1)(\mu_{\omega}-1.15\sigma_{\omega}),(m+1)(\mu_{\omega}+1.15\sigma_{\omega})\right].$
For the two lobes not to overlap, we need
$m(\mu_{\omega}+1.15\sigma_{\omega})<(m+1)(\mu_{\omega}-1.15\sigma_{\omega})$
which is equivalent to $\mu_{\omega}>1.15(2m+1)\sigma_{\omega}$ or $\mu_{v}>1.15(2m+1)\sigma_{v}$. This condition is reasonable to be met for harmonics $m\in\{1,2,3\}$ that are of primary interest.
The THC for the total DWPT load under S4 can be computed from (17). According to Lemma 4, the DWPT load has precisely the same THC under S2 and S4. This is because the Gaussian PDF integrates to unity as a function of $\omega$. This observation also implies that the energy contributed by the lobe centered around $m\mu_{\omega}$ under S4 equals the energy of harmonic $m\omega_{0}$ under S2. Interestingly, scenario S2 can be derived as the limiting case of S4 when $\sigma_{v}$ tends to zero and $\mu_{v}=v$. This originates from the fact that the Dirac delta function is the limiting case of the Gaussian PDF for vanishing variance.
To ease comparison with previous scenarios, consider again deterministic EV densities by setting $\sigma_{a}=0$ and $\mu_{a}=a$. Then, the DC term under S4 scales with $Nac_{0}$ and the Gaussian lobe centered around the $m$-th harmonic scales with $\sqrt{N}a|c_{m}|$. This is consistent because timings have been assumed independent across EVs as in S2.
@@
$\text{\emph{THC}}_{p}$ $=\text{\emph{THC}}_{h}$ $\text{for S${}_{1}$};$
$\text{\emph{THC}}_{p}$ $=\text{\emph{THC}}_{h}/\sqrt{N}$ $\text{for S${}_{2}$ and S${}_{4}$};and$
$\text{\emph{THC}}_{p}$ $=\text{\emph{THC}}_{h}\sqrt{Q/N}$ $\text{for S${}_{3}$}.$
Remark 3.
This work derives scaling laws for DWPT harmonics using the multiplicative model of DWPT load demand. In our later work [11], it is established that similar scaling laws apply to DWPT dynamics under the clipping model. This is because clipping only affects FS coefficients, whereas scaling laws are primarily determined by the EV timing distributions. Reference [11] also shows that the clipping model features lesser THC than the multiplicative model.
@@ V Periodograms of DWPT Load Demands
This question boils down to whether the periodogram of a finite-duration realization of $p(t)$ relates to its PSD. This is a well-studied topic in statistical signal processing [26]. Let us briefly review some key points. Consider a DWPT load realization $\hat{p}(t)$ spanning the finite interval $t\in[0,T_{p}]$ of duration $T_{p}$. Grid operators typically use windows of 1–5 min to compute spectra of signals in power system studies [21]. The periodogram of $\hat{p}(t)$ is computed as the squared magnitude of its Fourier transform as [26, Sec 4.3]
@@
In the three tests of this section, we computed numerically the periodograms of 1-min realizations $\hat{p}(t)$, generated according to S2S4. More specifically, we simulated the DWPT load consumed by an $L=16$ km (10 mile)-long ER segment following the specifications of the INDOT ER testbed. EV loads were computed as trapezoidal pulse trains. EV power densities $a_{n}$ were simulated as i.i.d. Gaussian random variables with mean $\mu_{a}=94.45$ kW/m and standard deviation $\sigma_{a}=5.55$ kW/m; see [6]. For S2 and S3, speed was set to $v=24.56$ m/s. At this speed, an EV takes about 10 min to cross the ER segment.
@@
Figure 4: Periodograms of 1-min DWPT loads simulated under scenario S2 for two values of $N$. The magnitude and location of harmonics agree with those predicted by Lemma 2. Moreover, as asserted by Lemma 2, the ratio of the two periodograms should be $10\log_{10}(192/64)^{2}=9.54$ dB at DC, and about $10\log_{10}\left({192}/{64}\right)=4.77$ dB at harmonic components.
@@
Periodogram under S2. To compute the periodogram under S2, EV timings were independently uniformly drawn from $[0,T)$, where $T=D/v$. Because independent timings are likely to occur under lighter traffic, we considered two relatively low values of the mean arrival rate $\lambda_{a}$, namely $0.1$ and $0.3$ EV/s, yielding $N=64$ and $192$ EVs, respectively. Figure 4 depicts the periodogram of a single 1-min realization of $\hat{p}(t)$ after applying a Hanning window. This periodogram differs from the PSD of Lemma 2 as it relies on a single, finite-duration realization $\hat{p}(t)$. Nonetheless, the locations of harmonic peaks coincide with those predicted by the analysis. For $N=64$ and $N=192$, the DC component of $\hat{p}(t)$ was $8.70$ MW and $26.13$ MW, respectively. Integrating the two periodograms around the first harmonic yielded $357$ kW and $680$ kW, respectively, for the two values of $N$. This verifies that although the DC component scales with $N$, harmonics scale with $\sqrt{N}$. In addition, we evaluated the THC for multiple realizations $\hat{p}(t)$ using (17) and found it to be 3–6% and 1–3% for the two values of $N$, indicating that the THC decreases with $N$.
@@
Figure 5: Periodograms of 1-min DWPT loads simulated under S2S4 for $N=192$ EVs. Harmonics increase in power under S3 due to platoons. The increase is about $10\log_{10}Q=6.98$ dB per Lemmas 23. The lobes under S4 are centered around the harmonics of S2 because $\mu_{v}=55$ mph. lobes reduce in magnitude and spread in bandwidth with $m$ per Lemma 4.
@@
Periodogram under S3. In this case, we simulated $N=192$ EVs moving in platoons of approximately $Q=5$ EVs per platoon. To capture imperfect synchronization, we first sampled the timing of each platoon independently uniformly from $[0,T)$. Then, for each EV within a platoon, we added an extra time shift drawn independently uniformly from $[0,T/5]$. Figure 5 illustrates the obtained periodogram. Although EVs within a platoon are not perfectly synchronized, the obtained spectrum contains higher harmonic content compared to the spectrum under S2. As dictated by Lemmas 2 and 3, although the DC components are identical (barring the deviations of a single periodogram from its mean), harmonics do increase in power under S3, as they now scale as $\sqrt{QN}$ instead of $\sqrt{N}$. In addition, we computed the THC under different realizations and found it to be around $3\%$ under imperfect synchronization and $16\%$ under perfect synchronization within platoons.